OpenStudy (anonymous):

Divide the following polynomial. (16x^2 -25) (4x + 5)

OpenStudy (abdullahm):

Hint: \(\sf\Large a^2-b^2 = (a+b)(a-b)\)

OpenStudy (abdullahm):

I'm assuming your question is \(\sf\LARGE \frac{16x^2-25}{4x+5}\)

OpenStudy (anonymous):

Yes!

OpenStudy (abdullahm):

simplify the numerator using the hint I gave you :)

OpenStudy (abdullahm):

if you need more help, look at the denominator and that should move you forward (:

OpenStudy (anonymous):

i still dont understand :(

OpenStudy (abdullahm):

what is \(\sf \sqrt{16} = ?\) \(\sf \sqrt{25} = ?\) \(\sf \sqrt{x^2}=?\) for this last one, hint is: \(\sf\Large\sqrt {x} = x ^{\frac{1}{2}}\)

OpenStudy (jdoe0001):

\(\large { 16x^2 -25\qquad \begin{cases} 16\to 4^2\\ 25\to 5^2 \end{cases}\implies \begin{array}{llll} 16x^2 -25\implies 4^2x^2-5^2\\ (4x)^2-5^2 \end{array} \\ \quad \\ \textit{keep in mind that } \begin{array}{llll} \textit{difference of squares} \\ (a-b)(a+b) = a^2-b^2\\ a^2-b^2 = (a-b)(a+b) \end{array} }\) so.. factor the numerator first, as aforementioned what factors do you get?